Ideal waist size:
For a woman with 40 inch hips the ideal waist is 32 inches. What is the maximum level if woman has 35 inch hips
Would you set the problem up like this:
40-35=5 then add 5 inches to the 32 inch waist, Right?
Total of 37 inches Or is it the 40-32=8 inches and add to 35inch waist?

Answer by nerdybill(7384)

For a woman with 40 inch hips the ideal waist is 32 inches. What is the maximum level if woman has 35 inch hips
.
No, it should be set up like a ratio problem.
Since the ideal size is "40 inch hips the ideal waist is 32 inches" the ratio is:
32:40
.
To calculate waist for 35 inch hips.
32:40 = 35:?
32/40 = 35/?
cross multiply to give:
32? = 40(35)
? = 40(35)/32
? = 43.75 inches